OpenStudy (anonymous):

What is the minimum product of two numbers whose difference is 48? What are the numbers? The minimum product is_____ The two number yield this product?______ Simplify your answer

OpenStudy (kinggeorge):

You have two numbers \(x, y\) such that \(x-y=48\) and you want to minimize \(xy\). First, rewrite the first equation to get \[x=48+y\]Then substitute that into the second equation. Thus, you get \[(y+48)y=y^2+48y\]Now we want to minimize that. Can you do that on your own?

OpenStudy (anonymous):

Need help please!

OpenStudy (kinggeorge):

You'll want to take the derivative of that function. So you get \(2y+48\). To find the minimum, find the zero of this function. This results in \[2y+48=0 \implies y=-\,{48 \over 2}=-24\]That's the value for y that gives you the minimum value. Now we need to find x. But wait! we have a function that solves for x in terms of y already. \[x=48+y\implies x=48-24=24\]So we have that \[x=24\]\[y=-24\] The product of these numbers is \[xy=24(-24)=-576\]